A small plant manufactures riding lawn mowers. The plant has fixed costs of GHC48,000 per day and a variable cost of GHC1,400 per unit produced. The mowers are sold for GHC1,800 each. The cost and revenue functions are in that order  and , where  is the total number of mowers produced and sold each day.How many units must be manufactured and sold each day for the company to break even?

1
Expert's answer
2021-03-09T02:20:15-0500

Break even is the point (x,y) where revenue = cost

Since y amounts are equal in cost and revenue equations

Let the,

Cost equations be:

y=48000+1400x

Revenue equation be:

y=1800x

Where x is the number of mowers produced and sold each day

Hence,

48000+1400x=1800x

x=48000/400

x=120

Substituting back in the revenue equation

y=1800*120

= 216,000